# Convergence Analysis and Acceleration of Fictitious Play for General Mean-Field Games via the Best Response

### v1

@@ 2.2.3 Non-potential MFGs
We first prove the decay property of $g$. A similar technique was used in [30] for anonymous games.
@@
Let the sequence $\{({\rho}^{(k)},{m}^{(k)})\},\{(\hat{\rho}^{(k)},\hat{m}^{(k)})\}$ be generated by the numerical scheme (16)-(17). If Assumption 2 holds and $f$ is monotone, then
@@
If in addition, Assumptions 1 holds, $r$ is directional $\lambda_{r}$-strongly convex between $({\rho}^{(k-1)},{m}^{(k-1)})$ and $(\hat{\rho}^{(k)},\hat{m}^{(k)})$, and between $(\hat{\rho}^{(k)},\hat{m}^{(k)})$ and $(\hat{\rho}^{(k+1)},\hat{m}^{(k+1)})$, then
@@
Therefore, by definition of gain of the best response, we have
@@
where $D=-\langle f(\rho^{(k-1)})-f(\rho^{(k)}),\hat{\rho}^{(k)}-\hat{\rho}^{(k+1)}\rangle_{Q_{T}}\geq 0$, by Lemma 2.10. And by Lemma 2.13, if Assumption 1 holds and $r$ is directional convex, then
@@
$A=$ $\underbrace{r(\rho^{(k)},m^{(k)})-r(\rho^{(k-1)},m^{(k-1)})+\langle f(\rho^{(k-1)}),{\rho}^{(k)}-\rho^{(k-1)}\rangle_{Q_{T}}+\langle f_{T},{\rho}^{(k)}_{T}-\rho^{(k-1)}_{T}\rangle_{\Omega}}_{A_{1}}$
@@
$A_{1}\leq$ $\delta_{k}(r(\hat{\rho}^{(k)},\hat{m}^{(k)})-r(\rho^{(k-1)},m^{(k-1)}))+\delta_{k}\langle f(\rho^{(k-1)}),\hat{\rho}^{(k)}-\rho^{(k-1)}\rangle_{Q_{T}}+\delta_{k}\langle f_{T},\hat{\rho}^{(k)}_{T}-\rho^{(k-1)}_{T}\rangle_{\Omega}$ (60)
@@
Remark 2.20 (Lack of Assumption 2).
If Assumption 2 does not hold, i.e. the terminal cost also depends on $\rho$, but $f$ and $f_{T}$ are monotone then under sufficient regularity and convexity assumptions, we have
@@
where $D=-\langle f(\rho^{(k-1)})-f(\rho^{(k)}),\hat{\rho}^{(k)}-\hat{\rho}^{(k+1)}\rangle_{Q_{T}}-\langle f_{T}(\rho^{(k-1)})-f_{T}(\rho^{(k)}),\hat{\rho}_{T}^{(k)}-\hat{\rho}_{T}^{(k+1)}\rangle_{\Omega}$. As long as $D<\delta_{k}g(\rho^{(k-1)},m^{(k-1)})$, we still have the decay of $g$.
@@
If $f$ is weakly monotone, i.e. there exists some $\mu_{f}\geq 0$ such that for any $\rho,\rho^{\prime}\in\mathcal{P}(\Omega)$,
$\int_{\Omega}(f(x,\rho)-f(x,\rho^{\prime}))\mathrm{d}(\rho-\rho^{\prime})(x)\geq-\mu_{f}\|\rho-\rho^{\prime}\|_{\Omega}^{2},$
@@
Let the sequence $\{({\rho}^{(k)},{m}^{(k)})\},\{(\hat{\rho}^{(k)},\hat{m}^{(k)})\}$ be generated by the numerical scheme (16)-(17). Under Assumption 1-2, if $f$ is monotone and $r$ is directional $\lambda_{r}$-strongly convex between $(\hat{\rho}^{(k)},\hat{m}^{(k)})$ and $(\hat{\rho}^{(k+1)},\hat{m}^{(k+1)})$, and between $({\rho}^{(k)},{m}^{(k)})$ and $(\hat{\rho}^{(k+1)},\hat{m}^{(k+1)})$, for any $k$, then we have
@@
Telescoping from $n=1$ to $K$ gives (66).
@@
Taking $\delta_{k}=\delta$ in (69) directly gives (68). ∎
@@
Let $(\rho^{*},m^{*})$ be the Nash Equilibrium and the sequence $\{({\rho}^{(k)},{m}^{(k)})\},\{(\hat{\rho}^{(k)},\hat{m}^{(k)})\}$ be generated by the numerical scheme (16)-(17). Under Assumptions 1-2, and assumptions of Theorem 2.22, if $f$ is $\lambda_{f}$ monotone, $r$ is directional $L_{r}$-Lipschitz smooth between $(\rho^{0},m^{0})$ and its best response $(\hat{\rho}^{1},\hat{m}^{1})$, and $r$ is directional $\lambda_{r}$-strongly convex between $(\rho^{0},m^{0})$ and $(\rho^{*},m^{*})$, then we have
$\|\rho^{(K)}-\rho^{*}\|_{Q_{T}}\leq\frac{L_{f}(L_{f}+\lambda_{f})}{\lambda_{f}\lambda_{r}}\sqrt{\frac{L_{r}}{\lambda_{r}}}\exp\left(\frac{1}{2}\sum_{k=1}^{K}(c\delta_{k}^{2}-\delta_{k})\right)\|\rho^{(0)}-\rho^{*}\|_{Q_{T}},$ (71)
@@
2.2.4 Acceleration: back-tracking line search
Theorem 2.22 demonstrates that the weight $\delta_{k}$ fundamentally determines the convergence rate. When constants $L_{f}$ and $\lambda_{r}$ are known, a fixed weight can be chosen to ensure linear convergence. However, in practice, these constants may be unknown, and $\lambda_{r}$ may vary with each iteration. In such cases, a common approach is to set $\delta_{k}=\frac{1}{k+1}$, as in [19] and other prior fictitious play studies, or to use the Frank-Wolfe weight $\delta_{k}=\frac{2}{k+2}$, which improves the convergence rate. However, both choices result in sublinear convergence that slows as $k$ increases. To address this, in this section, we adopt a backtracking line search strategy from optimization, allowing for a dynamic selection of an optimal weight $\delta_{k}$ to enhance the convergence rate. The numerical experiment in Section 4.3 illustrates the effectiveness of this acceleration.
Recall that the key to the convergence of fictitious play is Lemma 2.19, which states
@@ 2.2.4 Acceleration: back-tracking line search
and to ensure fast convergence, we need to maximize this improvement. When the best-response mapping is continuous, $D=\mathcal{O}(\delta_{k}^{2})$, and a sufficiently small $\delta_{k}$ guarantees positive improvement. Thus, decaying weights such as $\delta_{k}=\frac{1}{k+1}$ or $\delta_{k}=\frac{1}{k+2}$ will eventually lead to convergence. However, with these small values of $\delta_{k}$, the improvement in the gain function is only $\mathcal{O}(\delta_{k})$, which slows down convergence. Therefore, it is essential to choose the largest feasible $\delta_{k}$ that still ensures a positive improvement.
We adopt the backtracking line search strategy from optimization to achieve this. Algorithm 2 outlines the process, where $y\leftarrow z$ denotes assigning the value of $z$ to $y$. Each iteration within the while-loop requires solving the HJB and FP equations once. The backtracking line search aims to find the smallest non-negative integer $n$ such that $\delta_{k}=\delta\beta^{n}$ and $D\leq\zeta\delta_{k}g(\rho^{(k-1)},m^{(k-1)})$, where $\delta\in(0,1]$ and $\beta,\zeta\in(0,1)$ are preselected parameters. The parameter $\zeta$ controls the amount of gain improvement, specifically $(1-\zeta)\delta_{k}g(\rho^{(k-1)},m^{(k-1)})$. A smaller $\zeta$ yields greater improvement but may require more iterations to satisfy or may be difficult to achieve. The parameter $\beta$ determines the shrinkage rate for $\delta_{k}$ when the target improvement is unmet. A smaller $\beta$ results in fewer iterations to find a valid $\delta_{k}$ but may lead to overly modest improvements.
@@
According to Theorem 2.22, the weight (step-size) $\delta_{k}=\frac{\alpha}{k+\alpha},(\alpha>0)$ gives a convergence rate $\frac{1}{k^{\alpha}}$. But this does not suggest that larger $\alpha$ is in practice better because the constant also grows exponentially with $\alpha$. This is also noticed and discussed in conditional gradient method [45] and in potential games setting [41].
@@
[31] observes that fictitious play with $\delta_{k}=\frac{1}{k}$ converges very slowly as the number of iterations grows, primarily because excessive weight is given to inaccurate initial guesses and historical data. Without a theoretical explanation, [31] suggests and empirically demonstrates that when $({\rho}^{(k)},{m}^{(k)})$ is near equilibrium, restarting fictitious play with $({\rho}^{(k)},{m}^{(k)})$ keeps the value of $k$ small, thereby accelerating convergence. Our analysis reveals that this slow convergence for large $k$ stems from the sublinear convergence rate. Restarting the algorithm with a small $k$ results in a large $\delta_{k}$, which could prevent the decay property of the gain function $g$ from holding and lead to a large increase of $g$. Thus, the restart strategy primarily aids convergence asymptotically as $k$ increases.
@@
[41] also proposes a back-tracking line search strategy to determine the weight $\delta_{k}$. However, their strategy requires the existence of the potential and is only applicable to potential MFGs.
2.2.5 Potential MFGs
In the end, we discuss the convergence analysis for potential MFGs and highlight the difference to the non-potential case. We also remark on the connection between fictitious play and the Frank-Wolfe algorithm and compare fictitious play with other optimization-based algorithms on potential MFGs.
If the game is potential, we denote
$s(\rho):=\int_{0}^{T}F(\rho_{t})\mathrm{d}t,\quad s_{T}(\rho_{T}):=F_{T}(\rho_{T}),\quad j(\rho,m):=r(\rho,m)+s(\rho)+s_{T}(\rho_{T}).$
Under sufficient regularity assumptions, $(\rho^{*},m^{*})$ is the Nash Equilibrium if and only if
@@ 2.2.5 Potential MFGs
If we assume there exists such $(\rho^{*},m^{*})$, then we can define the primal gap functional
$h(\rho,m):=j(\rho,m)-j(\rho^{*},m^{*}),$
and $(\rho^{*},m^{*})$ being a Nash Equilibrium is equivalent to $h(\rho^{*},m^{*})=0$. Since for any $\rho,\tilde{\rho}$, it holds that
$\langle\nabla s(\rho),\tilde{\rho}\rangle_{Q_{T}}=\langle f(\rho),\tilde{\rho}\rangle_{Q_{T}},\quad\langle\nabla s_{T}(\rho_{T}),\tilde{\rho}_{T}\rangle_{\Omega}=\langle f_{T}(\rho_{T}),\tilde{\rho}_{T}\rangle_{\Omega},$
the fictitious play Algorithm 1 is equivalent to the scheme,
$(\hat{\rho}^{(k)},\hat{m}^{(k)})\in\operatorname*{argmin}_{(\tilde{\rho},\tilde{m})\in C}r(\tilde{\rho},\tilde{m})+\langle\nabla s(\rho^{(k-1)}),\tilde{\rho}\rangle_{Q_{T}}+\langle\nabla s_{T}(\rho_{T}^{(k-1)}),\tilde{\rho}_{T}\rangle_{\Omega},$ (72)
[41] notices that the scheme (72)-(73) is exactly applying the generalized conditional gradient to the variational formulation of potential MFGs (8). Generalized conditional gradient [12, 36] is a variant of the projection-free Frank-Wolfe algorithm [35]. According to [11], $\delta_{k}=\frac{2}{k+2}$ is called open loop step-size in Frank-Wolfe and is a standard choice when the information of the function is not available. The proof techniques in [41] adapt a lot from the techniques for the generalized conditional gradient in [36] and rely greatly on the gap function $h$, which is the primal gap in the literature on the Frank-Wolfe algorithm and generalized conditional gradient.
We note that when the game is potential, the gain of the best response $g$ is actually the dual gap,
$g(\rho,m)=\max_{\tilde{\rho},\tilde{m}}r(\rho,m)-r(\tilde{\rho},\tilde{m})+\langle\nabla s(\rho),\rho-\tilde{\rho}\rangle_{Q_{T}}+\langle\nabla s_{T}(\rho_{T}),\rho_{T}-\tilde{\rho}_{T}\rangle_{\Omega}.$
However, in the literature on the generalized conditional gradient, most studies on the dual gap are based on its relations with the primal gap, which non-potential MFG lacks. The study of the relation between the fixed point and gain of the best response (dual gap) in Section 2.2.2 of this paper is very important to adapt the convergence analysis from potential MFGs to non-potential MFGs.
In the rest of this section, we give a simplified convergence analysis for potential MFGs. Similarly, to prove the convergence for potential MFGs, the key is to establish the decay property of the primal gap, which does not require the directional strong convexity of $r$.
@@
Under Assumption 2, if $f$ has a convex potential $F$ and there exists $L_{f}>0$ such that $F$ is $L_{f}$-Lipschitz smooth, then the sequence generated by numerical scheme (72)-(73) satisfies
@@
By definition,
$h(\rho^{(k)},m^{(k)})-h(\rho^{(k-1)},m^{(k-1)})$ (75)
$=$ $j(\rho^{(k)},m^{(k)})-j(\rho^{(k-1)},m^{(k-1)})$
$=$ $r(\rho^{(k)},m^{(k)})-r(\rho^{(k-1)},m^{(k-1)})+s(\rho^{(k)})-s(\rho^{(k-1)})+\langle f_{T},\rho^{(k)}-\rho^{(k-1)}\rangle_{\Omega}.$
The convexity of $r$ and (73) give
$r(\rho^{(k)},m^{(k)})-r(\rho^{(k-1)},m^{(k-1)})\leq\delta_{k}(r(\hat{\rho}^{(k)},\hat{m}^{(k)})-r(\rho^{(k-1)},m^{(k-1)})).$ (76)
The $L_{f}$-smooth of $s$ and (73) give
$s(\rho^{(k)})-s(\rho^{(k-1)})\leq$ $\langle\nabla s(\rho^{(k-1)}),(\rho^{(k)}-\rho^{(k-1)})\rangle_{Q_{T}}+\frac{L_{f}}{2}\|\rho^{(k)}-\rho^{(k-1)}\|^{2}_{Q_{T}}$ (77)
$=$ $\delta_{k}\langle\nabla s(\rho^{(k-1)}),(\hat{\rho}^{(k)}-\rho^{(k-1)})\rangle_{Q_{T}}+\delta_{k}^{2}\frac{L_{f}}{2}\|\hat{\rho}^{(k)}-\rho^{(k-1)}\|^{2}_{Q_{T}}$
By combining (75), (76) and (77), we have
$h(\rho^{(k)},m^{(k)})-h(\rho^{(k-1)},m^{(k-1)})\leq$ $\delta_{k}\left(r(\hat{\rho}^{(k)},\hat{m}^{(k)})+\langle\nabla s(\rho^{(k-1)}),\hat{\rho}^{(k)}\rangle_{Q_{T}}+\langle f_{T},\hat{\rho}_{T}^{(k)}\rangle_{\Omega}\right)$ (78)
$-\delta_{k}\left(r(\rho^{(k-1)},m^{(k-1)})+\langle\nabla s(\rho^{(k-1)}),\rho^{(k-1)}\rangle_{Q_{T}}+\langle f_{T},\rho_{T}^{(k-1)}\rangle_{\Omega}\right)$
@@
The optimality of $(\hat{\rho}^{(k)},\hat{m}^{(k)})$ (72) gives
$r(\hat{\rho}^{(k)},\hat{m}^{(k)})+\langle\nabla s(\rho^{(k-1)}),\hat{\rho}^{(k)}\rangle_{Q_{T}}+\langle f_{T},\hat{\rho}_{T}^{(k)}\rangle_{\Omega}\leq r(\rho^{*},m^{*})+\langle\nabla s(\rho^{(k-1)}),\rho^{*}\rangle_{Q_{T}}+\langle f_{T},\rho_{T}^{*}\rangle_{\Omega}$ (79)
and the convexity of $s$ gives
$\langle\nabla s(\rho^{(k-1)}),\rho^{*}\rangle_{Q_{T}}-\langle\nabla s(\rho^{(k-1)}),\rho^{(k-1)}\rangle_{Q_{T}}\leq s(\rho^{*})-s(\rho^{(k-1)}).$ (80)

### v2

@@ 3.4 Non-potential MFGs
We first prove the decay property of $g$. A similar technique was used in hadikhanloo2017anonymous for anonymous games.
@@
Let the sequence $\{({\rho}^{(k)},{m}^{(k)})\},\{(\hat{\rho}^{(k)},\hat{m}^{(k)})\}$ be generated by the numerical scheme (24)-(25). If Assumption 3.6 holds and $f$ is monotone, then
@@
If in addition, Assumptions 3.5 holds, $R$ is directional $\lambda_{R}$-strongly convex between $({\rho}^{(k-1)},{m}^{(k-1)})$ and $(\hat{\rho}^{(k)},\hat{m}^{(k)})$, and between $(\hat{\rho}^{(k)},\hat{m}^{(k)})$ and $(\hat{\rho}^{(k+1)},\hat{m}^{(k+1)})$, then
@@
Therefore, we have
@@
where $D=-\langle f(\rho^{(k-1)})-f(\rho^{(k)}),\hat{\rho}^{(k)}-\hat{\rho}^{(k+1)}\rangle_{Q_{T}}\geq 0$, by Proposition 3.11. And by Lemma 3.14, if Assumption 3.5 holds and $R$ is directional convex, then
@@
$A=$ $\underbrace{R(\rho^{(k)},m^{(k)})-R(\rho^{(k-1)},m^{(k-1)})+\langle f(\rho^{(k-1)}),{\rho}^{(k)}-\rho^{(k-1)}\rangle_{Q_{T}}+\langle f_{T},{\rho}^{(k)}_{T}-\rho^{(k-1)}_{T}\rangle_{\mathbb{T}^{d}}}_{A_{1}}$
@@
$A_{1}\leq$ $\delta_{k}(R(\hat{\rho}^{(k)},\hat{m}^{(k)})-R(\rho^{(k-1)},m^{(k-1)}))+\delta_{k}\langle f(\rho^{(k-1)}),\hat{\rho}^{(k)}-\rho^{(k-1)}\rangle_{Q_{T}}+\delta_{k}\langle f_{T},\hat{\rho}^{(k)}_{T}-\rho^{(k-1)}_{T}\rangle_{\mathbb{T}^{d}}$ (69)
@@
Remark 3.17 (Lack of Assumption 3.6).
If Assumption 3.6 does not hold, i.e. the terminal cost also depends on $\rho$, but $f$ and $f_{T}$ are monotone, then under sufficient regularity and convexity assumptions, we have
@@
where
$D=-\langle f(\rho^{(k-1)})-f(\rho^{(k)}),\hat{\rho}^{(k)}-\hat{\rho}^{(k+1)}\rangle_{Q_{T}}-\langle f_{T}(\rho^{(k-1)})-f_{T}(\rho^{(k)}),\hat{\rho}_{T}^{(k)}-\hat{\rho}_{T}^{(k+1)}\rangle_{\mathbb{T}^{d}}.$
As long as $D<\delta_{k}g(\rho^{(k-1)},m^{(k-1)})$, we still have the decay of $g$.
@@
If $f$ is weakly monotone, i.e. there exists some $\mu_{f}\geq 0$ such that for any $\rho,\rho^{\prime}\in\mathcal{P}(\mathbb{T}^{d})\cap L^{2}(\mathbb{T}^{d})$,
$\int_{\mathbb{T}^{d}}(f(x,\rho)-f(x,\rho^{\prime}))\mathrm{d}(\rho-\rho^{\prime})(x)\geq-\mu_{f}\|\rho-\rho^{\prime}\|_{\mathbb{T}^{d}}^{2},$
@@
Let the sequence $\{({\rho}^{(k)},{m}^{(k)})\},\{(\hat{\rho}^{(k)},\hat{m}^{(k)})\}$ be generated by the numerical scheme (24)-(25). Under Assumption 3.5-3.6, if in addition $f$ is monotone and $R$ is directional $\lambda_{R}$-strongly convex between $(\hat{\rho}^{(k)},\hat{m}^{(k)})$ and $(\hat{\rho}^{(k+1)},\hat{m}^{(k+1)})$, and between $({\rho}^{(k)},{m}^{(k)})$ and $(\hat{\rho}^{(k+1)},\hat{m}^{(k+1)})$, for any $k$, then we have
@@
Telescoping from $k=1$ to $K$ gives (75).
@@
Taking $\delta_{k}=\delta$ in (79) and telescoping gives (77). ∎
@@
Let $(\rho^{*},m^{*})$ be the Nash Equilibrium and the sequence $\{({\rho}^{(k)},{m}^{(k)})\},\{(\hat{\rho}^{(k)},\hat{m}^{(k)})\}$ be generated by the numerical scheme (24)-(25). Under Assumptions 3.5-3.6, and assumptions of Theorem 3.19, if in addition $f$ is $\lambda_{f}$ monotone, then we have
$\|\rho^{(K)}-\rho^{*}\|_{Q_{T}}\leq\frac{L_{f}}{\lambda_{f}}\sqrt{\frac{2}{\lambda_{R}}g(\rho^{0},m^{0})}\exp\left(\frac{1}{2}\sum_{k=1}^{K}(c\delta_{k}^{2}-\delta_{k})\right),$ (80)
@@
3.5 Acceleration: back-tracking line search
Theorem 3.19 demonstrates that the weight $\delta_{k}$ fundamentally determines the convergence rate. Recall that the key to the convergence of fictitious play is Lemma 3.16, which states
@@ 3.5 Acceleration: back-tracking line search
and to ensure fast convergence, we need to maximize this improvement. When the best-response mapping is continuous, $D=\mathcal{O}(\delta_{k}^{2})$, and a sufficiently small $\delta_{k}$ guarantees positive improvement. Thus, decaying weights such as $\delta_{k}=\frac{1}{k+1}$ in cardaliaguet2017ficplay or $\delta_{k}=\frac{2}{k+2}$ will eventually lead to convergence. However, since the improvement in the gain function is $\mathcal{O}(\delta_{k})$, with these choices of $\delta_{k}$, the convergence slows down as $k$ increases.
Therefore, it is essential to choose the largest feasible $\delta_{k}$ that still ensures a positive improvement. In this section, we adopt a backtracking line search strategy from optimization, allowing for a dynamic selection of an optimal weight $\delta_{k}$ to enhance the convergence rate. The numerical experiment in Section 5.4 illustrates the effectiveness of this acceleration.
Algorithm 2 outlines the process, where $a\leftarrow b$ denotes assigning the value of $b$ to $a$. Each iteration within the while-loop requires solving the HJB and FP equations once. The backtracking line search aims to find the smallest non-negative integer $n$ such that $\delta_{k}=\delta\beta^{n}$ and $D\leq\zeta\delta_{k}g(\rho^{(k-1)},m^{(k-1)})$, where $\delta\in(0,1]$ and $\beta,\zeta\in(0,1)$ are preselected parameters. The parameter $\zeta$ controls the amount of gain improvement, specifically $(1-\zeta)\delta_{k}g(\rho^{(k-1)},m^{(k-1)})$. A smaller $\zeta$ yields greater improvement but may require more iterations to satisfy or may be difficult to achieve. The parameter $\beta$ determines the shrinkage rate for $\delta_{k}$ when the target improvement is unmet. A smaller $\beta$ results in fewer iterations to find a valid $\delta_{k}$ but may lead to overly modest improvements.
@@
According to Theorem 3.19, the weight (step-size) $\delta_{k}=\frac{\alpha}{k+\alpha},(\alpha>0)$ gives a convergence rate $\frac{1}{k^{\alpha}}$. But this does not suggest that a larger $\alpha$ is in practice better because the constant also grows exponentially with $\alpha$. This is also noticed and discussed in the conditional gradient method nesterov2018gcg-primaldual and in the potential games setting Lavigne2023gcg-ficplay .
@@
hadikhanloo2019finite observes that fictitious play with $\delta_{k}=\frac{1}{k}$ converges very slowly as the number of iterations grows, primarily because excessive weight is given to inaccurate initial guesses and historical data. Without a theoretical explanation, hadikhanloo2019finite suggests and empirically demonstrates that when $({\rho}^{(k)},{m}^{(k)})$ is near equilibrium, restarting fictitious play with $({\rho}^{(k)},{m}^{(k)})$ keeps the value of $k$ small, thereby accelerating convergence. Our analysis reveals that this slow convergence for large $k$ stems from the sublinear convergence rate. Restarting the algorithm with a small $k$ results in a large $\delta_{k}$, which could prevent the decay property of the gain function $g$ from holding and lead to a large increase of $g$. Thus, the restart strategy primarily aids convergence asymptotically as $k$ increases.
@@
Lavigne2023gcg-ficplay also proposes a back-tracking line search strategy to determine the weight $\delta_{k}$. However, their strategy requires the existence of the potential and is only applicable to potential MFGs.
3.6 Potential MFGs
Finally, we turn our attention to potential MFGs. We begin by examining the relationship between our Lyapunov function $g$ and the one used in Lavigne2023gcg-ficplay in the context of potential MFGs. Next, we present a simplified convergence proof tailored to the potential MFGs and highlight the key differences from the general, non-potential case. We conclude by comparing fictitious play with other optimization-based algorithms designed for potential MFGs.
Throughout Section 3.6, we assume that Assumptions 3.5 and 3.6 hold. In addition, we introduce the following assumption:
Assumption 3.24.
$f$ is derived from a functional $F:\mathcal{P}(\mathbb{T}^{d})\to\mathbb{R}$.
Assumption 3.6 further implies that $f_{T}$ is also derived from a potential of linear form
$F_{T}:\mathcal{P}(\mathbb{T}^{d})\to\mathbb{R},\quad\rho\mapsto\int_{\mathbb{T}^{d}}f_{T}(x)\,\mathrm{d}\rho(x).$
Under these conditions, and by Proposition 2.6, the fixed-point formulation of the MFG system (2) can be equivalently expressed as the convex optimization problem given in the variational formulation (26).
Lavigne2023gcg-ficplay points out that the fictitious play (24)-(25) is equivalent to applying the generalized conditional gradient (GCG) to the variational formulation (26). GCG Bredies2009gcg-itershrink ; Kunisch2024gcg-fastcvg is a projection-free optimization algorithm and can be viewed as a variant of the Frank-Wolfe method jaggi13gcg-revisiting-cg . Consider the optimization problem
$\min_{z\in Z}\quad j(z):=r(z)+s(z)$ (81)
where $Z\subset\mathbb{R}^{d}$ is a convex set, $r$ is possibly non-differentiable, and $s$ is differentiable. At iteration $k$, the GCG algorithm linearizes $s$ at $z^{(k-1)}$ and solves
$\hat{z}^{(k)}=\operatorname*{argmin}_{\tilde{z}\in Z}~{}r(\tilde{z})+\langle\nabla s(z^{(k-1)}),\tilde{z}\rangle,$ (82)
followed by the update
$z^{(k)}=(1-\delta_{k})z^{(k-1)}+\delta_{k}\hat{z}^{(k)}.$ (83)
Assuming $z^{*}\in Z$ is a minimizer, the primal gap is defined as
$h(z):=j(z)-j(z^{*})$ (84)
and the dual gap is defined as
$g(z):=\max_{\tilde{z}\in C}\{r(z)-r(\tilde{z})+\langle\nabla s(z),z-\tilde{z}\rangle\}=r(z)+\langle\nabla s(z),z\rangle-\min_{\tilde{z}\in C}\{r(\tilde{z})+\langle\nabla s(z),\tilde{z}\rangle\}$ (85)
In the MFG setting, at each iteration $(\rho^{(k-1)},m^{(k-1)})$, solving for the best response (24) corresponds to (82) by linearizing the interaction and terminal costs. And (25) corresponds to (83). Assume $(\rho^{*},m^{*})$ is a Nash Equilibrium with $\rho^{*}>0$ and satisfies
@@ 3.6 Potential MFGs
Since the objective $J$ is convex, the primal gap
$h(\rho,m):=J(\rho,m)-J(\rho^{*},m^{*})$ (86)
quantifies the suboptimality of a candidate pair $(\rho,m)$ relative to the Nash equilibrium. The dual gap in this context corresponds exactly to the gain of the best response $g$ defined in (23).
The proof techniques in Lavigne2023gcg-ficplay draw heavily from the analysis of the GCG method in Kunisch2024gcg-fastcvg , relying primarily on the decay of the primal gap $h$. Although the dual gap $g$ is also well-studied in the literature on Frank-Wolfe and GCG methods, its analysis is typically grounded in its relationship to the primal gap. In non-potential MFGs, where such a relationship may not exist, directly analyzing the dual gap becomes essential. To this end, Section 3.3 of this paper establishes a crucial connection between the fixed point of the best response mapping and the dual gap (i.e., the gain of the best response), providing the foundation for extending convergence analysis from potential to general MFGs.
Although our convergence analysis is formulated directly in terms of the dual gap $g$, we first show that in the potential setting, $g$ is equivalent to the primal gap $h$. This connection serves as a consistency check and helps unify the analysis across both frameworks.
Proposition 3.25.
Under Assumptions 3.5, 3.6 and 3.24, if $F$ is convex on $\mathcal{P}(\mathbb{T}^{d})$ under $L^{2}$ metric and $R$ is $\lambda_{R}$-strongly convex between $(\rho,m)$ and its best response, then
Proof.
We first show that $h(\rho,m)\leq g(\rho,m)$.
$h(\rho,m)$ $=R(\rho,m)-R(\rho^{*},m^{*})+\int_{0}^{T}F(\rho_{t})-F(\rho_{t}^{*})\mathrm{d}t+\langle f_{T},\rho_{T}-\rho^{*}_{T}\rangle_{\mathbb{T}^{d}}$ (88)
$\leq R(\rho,m)-R(\rho^{*},m^{*})+\langle f(\rho),\rho-\rho^{*}\rangle_{Q_{T}}+\langle f_{T},\rho_{T}-\rho^{*}_{T}\rangle_{\mathbb{T}^{d}}$
$\leq g(\rho,m).$
Here, the first and third lines are by definition and the second line is by convexity of $F$.
Next, we denote $(\hat{\rho},\hat{m})$ as the best response of $(\rho,m)\in C$ and
$(\rho^{+},m^{+})=(1-\delta)({\rho},{m})+\delta(\hat{\rho},\hat{m}),$
where $\delta\in(0,1]$. Then the convexity of $R$ and the average step give
$R(\rho^{+},m^{+})-R(\rho,m)\leq\delta(R(\hat{\rho},\hat{m})-R(\rho,m)).$ (89)
The Lipschitz continuity of $f$ (Assumption 3.5) and the average step give
$\int_{0}^{T}F(\rho^{+})-F(\rho)\mathrm{d}t\leq$ $\langle f(\rho),\rho^{+}-\rho\rangle_{Q_{T}}+\frac{L_{f}}{2}\|\rho^{+}-\rho\|^{2}_{Q_{T}}$ (90)
$=$ $\delta\langle f(\rho),\hat{\rho}-\rho\rangle_{Q_{T}}+\delta^{2}\frac{L_{f}}{2}\|\hat{\rho}-\rho\|^{2}_{Q_{T}}$
By combining (86), (89) and (90), we have
$h(\rho^{+},m^{+})-h(\rho,m)$ (91)
$\leq$ $\delta\Big{(}R(\hat{\rho},\hat{m})+\langle f(\rho),\hat{\rho}\rangle_{Q_{T}}+\langle f_{T},\hat{\rho}_{T}\rangle_{\mathbb{T}^{d}}\Big{)}-\delta\Big{(}R(\rho,m)+\langle f(\rho),\rho\rangle_{Q_{T}}+\langle f_{T},\rho_{T}\rangle_{\mathbb{T}^{d}}\Big{)}+\delta^{2}\frac{L_{f}}{2}\|\hat{\rho}-\rho\|^{2}_{Q_{T}}$
$=$ $-\delta g(\rho,m)+\delta^{2}\frac{L_{f}}{2}\|\hat{\rho}-\rho\|^{2}.$
Since $h\geq 0$ and by Lemma 3.12, we have
In the rest of this section, we give a simplified convergence analysis for potential MFGs. To prove the convergence for potential MFGs, we establish the decay property of the primal gap, which does not require the directional strong convexity of $R$.
@@
Under assumptions 3.5, 3.6 and 3.24, if $F$ is convex on $\mathcal{P}(\mathbb{T}^{d})$ under $L^{2}$ metric, then the sequence generated by fictitious play (24)-(25) satisfies
@@
Repeating the analysis in (89), (90) and (91) gives us
$h(\rho^{(k)},m^{(k)})-h(\rho^{(k-1)},m^{(k-1)})\leq$ $\delta_{k}\left(R(\hat{\rho}^{(k)},\hat{m}^{(k)})+\langle f(\rho^{(k-1)}),\hat{\rho}^{(k)}\rangle_{Q_{T}}+\langle f_{T},\hat{\rho}_{T}^{(k)}\rangle_{\mathbb{T}^{d}}\right)$ (94)
$-\delta_{k}\left(R(\rho^{(k-1)},m^{(k-1)})+\langle f(\rho^{(k-1)}),\rho^{(k-1)}\rangle_{Q_{T}}+\langle f_{T},\rho_{T}^{(k-1)}\rangle_{\mathbb{T}^{d}}\right)$
@@
The optimality of $(\hat{\rho}^{(k)},\hat{m}^{(k)})$ gives
$R(\hat{\rho}^{(k)},\hat{m}^{(k)})+\langle f(\rho^{(k-1)}),\hat{\rho}^{(k)}\rangle_{Q_{T}}+\langle f_{T},\hat{\rho}_{T}^{(k)}\rangle_{\mathbb{T}^{d}}\leq R(\rho^{*},m^{*})+\langle f(\rho^{(k-1)}),\rho^{*}\rangle_{Q_{T}}+\langle f_{T},\rho_{T}^{*}\rangle_{\mathbb{T}^{d}}$ (95)
and the convexity of $F$ gives
$\langle f(\rho^{(k-1)}),\rho^{*}-\rho^{(k-1)}\rangle_{Q_{T}}\leq\int_{0}^{T}F(\rho^{*}_{t})-F(\rho^{(k-1)}_{t})\mathrm{d}t.$ (96)